# STARS: Self-supervised Tuning for 3D Action Recognition in Skeleton Sequences

### v1

@@ Abstract
Self-supervised pretraining methods with masked prediction demonstrate remarkable within-dataset performance in skeleton-based action recognition. However, we show that, unlike contrastive learning approaches, they do not produce well-separated clusters. Additionally, these methods struggle with generalization in few-shot settings. To address these issues, we propose Self-supervised Tuning for 3D Action Recognition in Skeleton sequences (STARS). Specifically, STARS first uses a masked prediction stage using an encoder-decoder architecture. It then employs nearest-neighbor contrastive learning to partially tune the weights of the encoder, enhancing the formation of semantic clusters for different actions. By tuning the encoder for a few epochs, and without using hand-crafted data augmentations, STARS achieves state-of-the-art self-supervised results in various benchmarks, including NTU-60, NTU-120, and PKU-MMD. In addition, STARS exhibits significantly better results than masked prediction models in few-shot settings, where the model has not seen the actions throughout pretraining. Project page: https://soroushmehraban.github.io/stars/
@@ 1 Introduction
Despite the significant efforts that have been made, how to learn a more discriminative representation of skeletons is still an issue for skeleton-based action recognition. We believe that integrating MAE-based approaches with contrastive-learning methods can enhance the generalizability of representations, while preserving the strong performance of MAE models in within-dataset evaluations. To this end, we propose Self-supervised Tuning for 3D Action Recognition in Skeleton Sequences (STARS), a simple and efficient self-supervised framework for 3D action representation learning. It is a sequential approach that initially uses MAE as the pretext task. In the subsequent stage, it trains a contrastive head in addition to partially tuning the encoder for a few epochs, motivating the representation to form distinct and well-separated clusters. Fig. 1 shows that STARS requires significantly less resources during pretraining compared to contrastive learning approaches. In addition, STARS outperforms both MAE and contrastive learning approaches. In summary, our main contributions are as follows:
@@ 4.2 Experimental Setup
Data Preprocessing: From an initial skeleton sequence, a consecutive segment is randomly trimmed with a proportion , where is sampled from the range [0.5, 1] during training and, similar to [28], remains fixed at 0.9 during testing. Subsequently, the segment is resized to a consistent length using bilinear interpolation. By default, is set to 120.
@@
Pre-training: The first stage follows the same setting as MAMP [28]. For the second stage, we use the AdamW optimizer with weight decay 0.01, betas (0.9, 0.95), and learning rate 0.001. In the second stage, we train the projection and prediction modules in addition to finetuning the encoder for 20 epochs. We employ layer-wise learning rate decay with a decay rate of 0.20. All the pretraining experiments are conducted using PyTorch on four NVIDIA A40 GPUs with a batch size of 32 per GPU.
4.3 Evaluation and Comparison
Method | Input | NTU-60 | NTU-120 | PKU-II | ||
XSub(%) | XView(%) | XSub(%) | XSet(%) | XSub(%) | ||
Other pretext tasks: | ||||||
LongTGAN [49] | Single-stream | 39.1 | 48.1 | - | - | 26.0 |
Contrastive Learning: | ||||||
ISC [39] | Single-stream | 76.3 | 85.2 | 67.1 | 67.9 | 36.0 |
CrosSCLR [23] | Three-stream | 77.8 | 83.4 | 67.9 | 66.7 | 21.2 |
AimCLR [16] | Three-stream | 78.9 | 83.8 | 68.2 | 68.8 | 39.5 |
CPM [47] | Single-stream | 78.7 | 84.9 | 68.7 | 69.6 | - |
PSTL [51] | Three-stream | 79.1 | 83.8 | 69.2 | 70.3 | 52.3 |
CMD [30] | Single-stream | 79.4 | 86.9 | 70.3 | 71.5 | 43.0 |
HaLP [34] | Single-stream | 79.7 | 86.8 | 71.1 | 72.2 | 43.5 |
HiCLR [48] | Three-stream | 80.4 | 85.5 | 70.0 | 70.4 | - |
HiCo-Transformer [8] | Single-stream | 81.1 | 88.6 | 72.8 | 74.1 | 49.4 |
SkeAttnCLR [18] | Three-stream | 82.0 | 86.5 | 77.1 | 80.0 | 55.5 |
I2MD [29] | Three-stream | 83.4 | 88.0 | 73.1 | 74.1 | 49.0 |
ActCLR [25] | Three-stream | 84.3 | 88.8 | 74.3 | 75.7 | - |
Masked Prediction: | ||||||
SkeletonMAE [43] | Single-stream | 74.8 | 77.7 | 72.5 | 73.5 | 36.1 |
MAMP [28] | Single-stream | 84.9 | 89.1 | 78.6 | 79.1 | 52.0* |
Masked Prediction + Contrastive Learning: | ||||||
STARS-3stage (Ours) | Single-stream | 86.3 | 90.7 | 79.3 | 80.6 | 52.2 |
STARS (Ours) | Single-stream | 87.1 | 90.9 | 79.9 | 80.8 | 52.7 |
@@ 4.3 Evaluation and Comparison
Colorization [46] | DGCNN | Three-stream | 88.0 | 94.9 | - | - |
Hi-TRS [6] | Transformer | Three-stream | 90.0 | 95.7 | 85.3 | 87.4 |
@@
Few-shot Evaluation Protocol: This protocol evaluates the scenario where only a small number of samples are labeled in the target dataset. This is crucial in practical applications like education, sports, and healthcare, where actions may not be clearly defined in publicly available datasets. In this protocol, we pretrain the model on NTU-60 (XSub) and evaluate it on the evaluation set of 60 novel actions on NTU-120 (XSub) using labeled sequences for each class in -shot setting. For the evaluation, we calculate the cosine distance between the test sequences and the exemplars, and use -nearest neighbors to determine the action. Tab. 5 compares different methods in the few-shot settings. Notably, MAMP demonstrates poor generalization performance, in contrast to its robust performance in transfer learning and evaluations within the dataset. By applying contrastive tuning, STARS surpasses contrastive learning approaches in all settings, demonstrating its strength in various evaluations.
Method | To PKU-II | |
NTU 60 | NTU 120 | |
LongTGAN [49] | 44.8 | - |
MS2L [24] | 45.8 | - |
ISC [39] | 51.1 | 52.3 |
CMD [30] | 56.0 | 57.0 |
HaLP+CMD [34] | 56.6 | 57.3 |
SkeletonMAE [43] | 58.4 | 61.0 |
MAMP [28] | 70.6 | 73.2 |
STARS-3stage (Ours) | 71.8 | 72.7 |
STARS (Ours) | 71.9 | 72.2 |
@@ 4.4 Ablation Study
Queue size: Fig. 3 (b), shows that queue size has minimal influence in pretraining, and 8k is used for all of our evaluations.
@@ 5 Conclusion
In this work, we proposed a sequential contrastive tuning method. We find that masked prediction methods, despite showing promising results in various within-dataset evaluations, cannot outperform contrastive learning based methods in few-shot settings. By using our STARS framework, we show that we can further enhance the masked prediction baseline while achieving competitive results in few-shot settings, outperforming other models in 5-shot setting. However, when the dataset size is limited for pretraining and for evaluations when encoder is fine-tuned with supervision, STARS cannot add significant value to its baseline.
@@
An alternative pretraining method is to do it in 3 stages. Figure 5 shows the 3-stage design. Initially, when we initialize the projector and predictor modules, we freeze the encoder weights. In the second stage, we exclusively train the projector and predictor modules until they can effectively differentiate between different sequences using the NNCLR approach. Finally, in the third stage, we fine-tune the encoder weights using layer-wise learning rate decay.
@@ Appendix B Semi-supervised Evaluation Results
Hyperparameter | Value |
Learning rate | 0.001 |
Batch size | 32 |
Augmentations | Mirroring & Rotation |
Queue size | 32,768 |
Momentum | 0.999 |
Temperature | 0.07 |
@@ Appendix C Ablation Hyper-parameters
Tab. 10 shows the hyperparameters used in DINOTuning strategy. For simplicity and because of limitation in resources, we used only two global views and did not use any local views in DINO. As shown in Tab. 11, we can see that incorporating local views led to a small improvement in performance. However, it came at the cost of significantly more resources. To be specific, we introduced two local views that randomly trimmed a section of the sequence between 40% and 80% and fed it only to the student network. With these additional views, we had to reduce the batch size to 16 and double the training time. Consequently, in our other experiments, we stuck to using only global views. We also used Sinkhorn-Knopp centering [3] a KoLeo regularizer [33] to help the convergence. Tab. 8 shows the hyperparameters used in MoCoTuning. Similar to previous approaches [16, 23], we used 32K as the queue size, 0.999 for the momentum and 0.07 for the temperature.
@@ C.1 Algorithm psuedo code
Algorithm 1 demonstrates the process in PyTorch-style pseudo code.
Appendix D Confusion Matrix

### v2

@@ Abstract
Self-supervised pretraining methods with masked prediction demonstrate remarkable within-dataset performance in skeleton-based action recognition. However, we show that, unlike contrastive learning approaches, they do not produce well-separated clusters. Additionally, these methods struggle with generalization in few-shot settings. To address these issues, we propose Self-supervised Tuning for 3D Action Recognition in Skeleton sequences (STARS). Specifically, STARS first uses a masked prediction stage using an encoder-decoder architecture. It then employs nearest-neighbor contrastive learning to partially tune the weights of the encoder, enhancing the formation of semantic clusters for different actions. By tuning the encoder for a few epochs, and without using hand-crafted data augmentations, STARS achieves state-of-the-art self-supervised results in various benchmarks, including NTU-60, NTU-120, and PKU-MMD. In addition, STARS exhibits significantly better results than masked prediction models in few-shot settings, where the model has not seen the actions throughout pretraining.
@@ 1 Introduction
Despite the significant efforts that have been made, how to learn a more discriminative representation of skeletons is still an issue for skeleton-based action recognition. MAE methods emphasize high-frequency details, while contrastive learning (CL) focuses on global semantics and low-frequency information [35]. Combining them can improve generalizability while preserving the strong within-dataset performance of MAE models. To this end, we propose Self-supervised Tuning for 3D Action Recognition in Skeleton Sequences (STARS), a simple and efficient self-supervised framework for 3D action representation learning. It is a sequential approach that initially uses MAE as the pretext task. In the subsequent stage, it trains a contrastive head in addition to partially tuning the encoder for a few epochs, motivating the representation to form distinct and well-separated clusters. Fig. 1 shows that STARS requires significantly less resources during pretraining compared to contrastive learning approaches. In addition, STARS outperforms both MAE and contrastive learning approaches.
In summary, our main contributions are as follows:
@@ 4.2 Experimental Setup
Data Preprocessing:
From an initial skeleton sequence, a consecutive segment is randomly trimmed with a proportion , where is sampled from the range [0.5, 1] during training and, similar to [31], remains fixed at 0.9 during testing. Subsequently, the segment is resized to a consistent length using bilinear interpolation. By default, is set to 120.
| Method | Input | NTU-60 | NTU-120 | PKU-I | PKU-II | ||
| XSub(%) | XView(%) | XSub(%) | XSet(%) | XSub(%) | XSub(%) | ||
| Other pretext tasks: | |||||||
| LongTGAN [54] | Single-stream | 39.1 | 48.1 | - | - | 67.7 | 26.0 |
| Contrastive Learning: | |||||||
| ISC [43] | Single-stream | 76.3 | 85.2 | 67.1 | 67.9 | 82.8 | 36.0 |
| CrosSCLR [25] | Three-stream | 77.8 | 83.4 | 67.9 | 66.7 | 84.9 | 21.2 |
| AimCLR [18] | Three-stream | 78.9 | 83.8 | 68.2 | 68.8 | 87.4 | 39.5 |
| CPM [51] | Single-stream | 78.7 | 84.9 | 68.7 | 69.6 | 88.8 | 48.3 |
| PSTL [56] | Three-stream | 79.1 | 83.8 | 69.2 | 70.3 | 89.2 | 52.3 |
| CMD [30] | Single-stream | 79.4 | 86.9 | 70.3 | 71.5 | - | 43.0 |
| HaLP [37] | Single-stream | 79.7 | 86.8 | 71.1 | 72.2 | - | 43.5 |
| HiCLR [52] | Three-stream | 80.4 | 85.5 | 70.0 | 70.4 | - | - |
| HiCo-Transformer [10] | Single-stream | 81.1 | 88.6 | 72.8 | 74.1 | 89.3 | 49.4 |
| SkeAttnCLR [20] | Three-stream | 82.0 | 86.5 | 77.1 | 80.0 | 89.5 | 55.5 |
| I2MD [32] | Three-stream | 83.4 | 88.0 | 73.1 | 74.1 | - | 49.0 |
| ActCLR [27] | Three-stream | 84.3 | 88.8 | 74.3 | 75.7 | - | - |
| UmURL [41] | Single-stream | 82.3 | 89.8 | 73.5 | 74.3 | - | 52.1 |
| Masked Prediction: | |||||||
| SkeletonMAE [47] | Single-stream | 74.8 | 77.7 | 72.5 | 73.5 | 82.8 | 36.1 |
| MAMP [31] | Single-stream | 84.9 | 89.1 | 78.6 | 79.1 | 91.9* | 52.0* |
| S-JEPA [1] | Single-stream | 85.3 | 89.8 | 79.6 | 79.9 | 92.2 | 53.5 |
| Masked Prediction + Contrastive Learning: | |||||||
| PCM3 [53] | Single-stream | 83.9 | 90.4 | 76.5 | 77.5 | - | 51.5 |
| STARS-3stage (Ours) | Single-stream | 86.3 | 90.7 | 79.3 | 80.6 | 91.5 | 52.2 |
| STARS (Ours) | Single-stream | 87.1 | 90.9 | 79.9 | 80.8 | 91.2 | 52.7 |
@@
Pre-training: The first stage follows the same setting as MAMP [31]. For the second stage, we use the AdamW optimizer with weight decay 0.01, betas (0.9, 0.95), and learning rate 0.001. In the second stage, we train the projection and prediction modules in addition to finetuning the encoder for 20 epochs, and the best representation is chosen based on K-NN (K=10) on validation data. We employ layer-wise learning rate decay with a decay rate of 0.20. All the pretraining experiments are conducted using PyTorch on four NVIDIA A40 GPUs with a batch size of 32 per GPU.
4.3 Evaluation and Comparison
@@
| Colorization [50] | Three-stream | DGCNN | 88.0 | 94.9 | - | - |
| Hi-TRS [8] | Three-stream | Transformer | 90.0 | 95.7 | 85.3 | 87.4 |
@@
Few-shot Evaluation Protocol: This protocol evaluates the scenario where only a small number of samples are labeled in the target dataset. This is crucial in practical applications like education, sports, and healthcare, where actions may not be clearly defined in publicly available datasets. In this protocol, we pretrain the model on NTU-60 (XSub) and evaluate it on the evaluation set of 60 novel actions on NTU-120 (XSub) using labeled sequences for each class in -shot setting. For the evaluation, we follow MotionBERT [57] and calculate the cosine distance between the test sequences and the exemplars, and use -nearest neighbors to determine the action. Tab. 5 compares different methods in the few-shot settings. Notably, MAMP demonstrates poor generalization performance, in contrast to its robust performance in transfer learning and evaluations within the dataset. By applying contrastive tuning, STARS surpasses contrastive learning approaches in all settings, demonstrating its strength in various evaluations.
| Method | To PKU-II | ||
| NTU 60 | NTU 120 | PKU-I | |
| LongTGAN [54] | 44.8 | - | 43.6 |
| MS2L [26] | 45.8 | - | 44.1 |
| ISC [43] | 51.1 | 52.3 | 45.1 |
| CMD [30] | 56.0 | 57.0 | - |
| HaLP+CMD [37] | 56.6 | 57.3 | - |
| SkeletonMAE [47] | 58.4 | 61.0 | 62.5 |
| MAMP [31] | 70.6 | 73.2 | 67.8* |
| STARS-3stage (Ours) | 71.8 | 72.7 | 67.6 |
| STARS (Ours) | 71.9 | 72.2 | 68.5 |
@@ 4.4 Ablation Study
Queue size: Fig. 4 (b) explores how different queue sizes affect model accuracy during contrastive tuning, evaluated using the KNN protocol. The results indicate that the queue size has little impact on performance during pretraining. Based on these findings, we chose a queue size of 8k for all our evaluations.
@@ 5 Conclusion
In this work, we proposed a sequential contrastive tuning method. We find that masked prediction methods, despite showing promising results in various within-dataset evaluations, cannot outperform contrastive learning based methods in few-shot settings. By using our STARS framework, we show that we can further enhance the masked prediction baseline while achieving competitive results in few-shot settings, outperforming other models in 5-shot setting. However, when the dataset size is limited for pretraining, and for evaluations when encoder is fine-tuned with supervision, STARS cannot add significant value to its baseline.
Supplementary Material
@@
An alternative pretraining method is to follow MAE-CT [24] and tune the encoder in 3 stages. Figure 5 shows the 3-stage design. Initially, when we initialize the projector and predictor modules, we freeze the encoder weights. In the second stage, we exclusively train the projector and predictor modules until they can effectively differentiate between different sequences using the NNCLR approach. Finally, in the third stage, we fine-tune the encoder weights using layer-wise learning rate decay. One motivation for this staged approach is the idea that the NNCLR head’s random weights could interfere with the representation quality by mapping the features into a random space, disrupting the learned structure. However, our findings challenge this assumption. The t-SNE visualization in Fig. 6 demonstrates that even with random NNCLR head weights, the cluster structure in the encoder’s output space remains intact in the new representation space. Furthermore, we observed with STARS-3stage that replicating this three-stage process not only increases training time but also leads to a drop in final accuracy. As a result, we use a two-stage design in our proposed STARS method.
@@ Appendix C Ablation Hyper-parameters
Tab. 9 shows the hyperparameters used in DINOTuning strategy. For simplicity and because of limitation in resources, we used only two global views and did not use any local views in DINO. As shown in Tab. 10, we can see that incorporating local views led to a small improvement in performance. However, it came at the cost of significantly more resources. To be specific, we introduced two local views that randomly trimmed a section of the sequence between 40% and 80% and fed it only to the student network. With these additional views, we had to reduce the batch size to 16 and double the training time. Consequently, in our other experiments, we stuck to using only global views. We also used Sinkhorn-Knopp centering [5] a KoLeo regularizer [36] to help the convergence. Tab 13 shows the hyperparameters used in MoCoTuning. Similar to previous approaches [18, 25], we used 32K as the queue size, 0.999 for the momentum and 0.07 for the temperature.
@@ C.1 Algorithm psuedo code
Algo. 1 demonstrates the process in PyTorch-style pseudo code.
Appendix D Additional Ablation Studies
| Method | NTU-60 | |
| XSub | XView | |
| MAMP (Baseline) | 63.1 | 80.3 |
| STARS | 79.9 | 88.6 |
| MAMP + NNCLR | 74.6 | 86.5 |
| K | NTU-60 | |
| XSub | XView | |
| 1 | 37.6 | 30.5 |
| 2 | 35.8 | 28.7 |
| 5 | 39.9 | 32.3 |
| 10 | 41.2 | 33.6 |
| Hyperparameter | Value |
| Learning rate | 0.001 |
| Batch size | 32 |
| Augmentations | Mirroring & Rotation |
| Queue size | 32,768 |
| Momentum | 0.999 |
| Temperature | 0.07 |
| Method | NTU-60 | |
| XSub | XView | |
| MAMP | 63.1 | 80.3 |
| STARS | 79.9 | 88.6 |
| MAE | 44.1 | 43.7 |
| STARS-MAE | 53.5 | 55.2 |
| Method | 1-shot | 2-shot | 5-shot |
| MAMP | 47.6 | 44.4 | 48.4 |
| STARS | 63.5 | 62.2 | 65.7 |
| MAE | 35.0 | 31.8 | 34.2 |
| STARS-MAE | 41.5 | 37.9 | 40.6 |
Combining MAMP and NNCLR.: One idea to tune the encoder in second stage is to combine NNCLR with MAMP and use both. Tab 11 compares this tuning stage with the proposed STARS. By including MAMP in the second stage of tuning, although it improves the final representation of encoder compared to the baseline (MAMP), it cannot perform as effective as STARS.
Training NNCLR from scratch: Table 12 presents the K-NN evaluation results for training the transformer from scratch using the NNCLR method for 300 epochs. Without augmentation, the model struggles to select positive samples in contrastive learning that truly represent the same actions. This happens because the encoder starts with random weights, which lack meaningful cluster separation. Consequently, the encoder fails to fully use the advantages of NNCLR in the second stage of STARS, resulting in poorer performance.
Using naive MAE instead of MAMP: Tab. 14 and Tab.15 present a comparison of K-NN and few-shot evaluations for a variant that uses naive MAE instead of MAMP in the first stage. While tuning in the second stage leads to significant improvements, the overall performance remains lower because MAE performs worse than MAMP in the initial stage.
| Batch size | K-NN Acc. (%) |
| 16 | 76.7 |
| 32 | 78.9 |
| 64 | 78.3 |
| 128 | 79.9 |
Effect of Batch Size. Table 16 reports the effect of varying the effective batch size on K-NN classification accuracy. Increasing the batch size, which corresponds to sampling more negatives, generally improves accuracy, with the best performance observed at a batch size of 128. This highlights the importance of large negative sets for robust representation.
| Setting | K-NN Acc. (%) |
| 74.3 | |
| 83.7 |
Effect of layer decay. As shown in Fig. 4 of the main paper, applying layer decay with yields the best performance. For comparison, we also evaluate the case of , where all layers are tuned uniformly. Table 17 compares the effect of different tuning strengths on -NN accuracy with . While full tuning () achieves moderate performance, reducing the tuning strength to substantially improves accuracy, suggesting that partial tuning helps preserve more generalizable representations.
| Method | Parameters (M) | MACs (G) |
| MAMP | 10.34 | 250.4 |
| STARS | 8.43 | 606.1 |
Parameters and MACs. Tab. 18 shows that STARS requires fewer parameters than MAMP, as it employs an MLP-based decoder instead of a transformer decoder. However, its MACs are significantly higher because training involves two forward passes for the contrastive loss as well as the additional computation needed to identify nearest neighbors and apply the loss.
Appendix E Confusion Matrix